# Benchmark Meta2 Evaluation - How to get it right

### v1

@@ 6 BAT uses in Related Work
Some works survey and analyze a field by utilizing BAT techniques. Liu2021DoQA check agreement across many QA datasets and conclude that since agreement is high, there is no need for more QA datasets. sun2023validity use correlations to show that Compositionality Benchmarks do not agree amongst themselves. They used Kendall-Tau and set as the high agreement threshold. Other works performed general efficient evaluation research and utilized BAT (Prabhu2024LifelongBE; Perlitz2023EfficientB; Polo2024tinyBenchmarksEL; Viswanathan2023Prompt2ModelGD). All of these works performed a thoughtful evaluation and large (reliable) rank correlation over all the models in the benchmarks. However, they did not consider the high correlations achieved in such settings (§3.2).
Other work relies on BAT to compare to a specific benchmark. Feng2024SampleEfficientHE automatically sample a small set of instructions as an efficient LLM benchmark, reducing human labor significantly. They show this still agrees with existing benchmarks. Similarly, Lei2023S3EvalAS and Viswanathan2023Prompt2ModelGD both propose a synthetic benchmark as a proxy and show good agreement with the original benchmark, although they differ in their methodology. Chang2023DoLM propose two benchmarks and use agreement to show that they capture the same phenomenon, and Mizrahi2023StateOW test agreement within the same benchmarks using different prompts. Li2024TreeEvalBE validate a new benchmark with 6 models of 3 sizes 7B,13B,33B with agreement alpaca(v2 )(alpaca_eval).
Last, Yuan2024PRobELMPR and (waldis2024holmes) show divergent validity by comparing their benchmark to established ones, showing low BAT scores.
@@ 7 Discussion and Conclusions
[…]
@@
It is not trivial how one should treat two benchmarks that are in high agreement with each other. If one is more convenient to run (e.g., doesn’t require costly metrics), then from a practical perspective, a user can simply choose it over the more expensive one. However, practitioners and researchers must not confuse high agreement with the notion that the benchmarks actually measure the exact same qualities. Among other things, this could lead to an erroneous view that new benchmarks are no longer needed, impeding new benchmark development. The community must also discriminate between correlations of model abilities (strong models are strong at many tasks) and correlations of the benchmarks themselves (the benchmarks actually measure the same qualities).
@@
Another point concerns the role of BAT for benchmark retirement, i.e., at what point do we decide that an old benchmark is no longer relevant and should be discarded. Currently the issue of retirement is viewed mainly from the perspective of saturation, where the community stops using benchmarks on which all new models succeed. However, another reason to retire benchmarks may be that the mixture of abilities models are expected to possess has shifted over time. In this scenario, BAT can reveal that a certain benchmark is no longer viable.
@@ 8 Limitations
We note that finding low agreement may mean two things, while both are negative, they should be addressed or interpreted differently. One option is that the benchmark measures something different from what it is supposed to and is hence not valid. That is the more common interpretation and calls for changes. Another option might be that the benchmark is just not reliable, intuitively its ranking is unstable and did not converge. In such cases, even the same benchmark may not agree with itself given small changes (subsets, seeds etc.), this usually calls for evaluating on more examples (choshen-etal-2024-navigating). There is a positive note to the same story, if a benchmark already shows a strong BAT in fine-grained evaluation (e.g., 5 models close to each other), it also means that it is quite reliable.
@@
In general, BAT needs a reference benchmark, or ideally multiple ones giving diverse measurements of the same construct. Still, choosing the right reference benchmarks might be tricky, and the results might be sensitive to this choice.
9 Appendices
9.1 Benchmarks used
The AGI Eval (Zhong2023AGIEvalAH) benchmark assesses models on human-level cognition and problem-solving tasks, which tests the real-world applicability of model outputs. Similarly, Alpaca (v2) (alpaca_eval) and its length-adjusted version (Dubois2024LengthControlledAA) focus on a model’s ability to follow complex instructions with the latter specifically addressing biases associated with output length.
HumanEval (Chen2021EvaluatingLL) presents code generation challenges, evaluating the syntactic correctness and logical soundness of model-generated code. Alongside, the HuggingFace OpenLLM Leaderboard (open-llm-leaderboard) employs the Eleuther AI Evaluation Harness (eval-harness) to test models on several key benchmarks such as ARC (clark2018think), HellaSwag (zellers2019hellaswag), MMLU (hendrycks2021measuring), TruthfulQA (lin2022truthfulqa), Winogrande (sakaguchi2021winogrande), and GSM8k (cobbe2021gsm8k). EQ-Bench (v2) (paech2023eqbench), measures the emotional intelligence of models, essential for applications that involve nuanced human interactions.
The MAGI (MAGIbenchmark) benchmark integrates challenging elements from MMLU and AGIEval to test complex reasoning and problem-solving capabilities of models. It is particularly effective in highlighting subtle performance differences among top-tier models. MMLU (Hendrycks2020MeasuringMM) assesses both general and specialized knowledge across various domains, providing a broad evaluation spectrum.
Further, benchmarks like Chatbot-Arena and MTBench (Zheng2023JudgingLW) focus on multi-turn conversation abilities, crucial for applications in customer service and virtual assistance. Lastly, Big Bench Hard (suzgun2022challenging) challenges models with complex text understanding and generation, pushing the limits of what natural language processing technologies can achieve. It is worth noting, that the HELM benchmark (liang2023holistic) was excluded from our analysis because there were few overlapping models with the other benchmarks.
9.2 Model Tier
Building on the importance of model proximity, another crucial factor in benchmark agreement is the tier of models being assessed. Current BAT practices often treat benchmarks as a uniform slab, disregarding the variations across different tiers of model performance. However, agreement might not be uniform across these tiers, and understanding this variance can provide deeper insights into benchmark reliability and model performance.
In Figure 7, we show that model tier significantly impacts benchmark agreement. Bottom-tier models exhibit higher agreement among themselves, with Kendall correlation coefficients just below 0.5. In contrast, middle-tier models show low agreement (coefficients below 0.2), and top-tier models demonstrate low to medium agreement (around 0.3).
One potential explanation for this phenomenon is the (lack of) reliability of the benchmark, as discussed in the introduction and literature (Perlitz2023EfficientB). Figure LABEL:fig:top_middle_bottom_std highlights that the standard deviation of scores buttom-ranked models is significantly higher than the rest. This might mean that there is some effect the goes beyond granularity or density, with older models being easier to differentiate (and gaining higher correlations to the models). However middle and top ranked models do not show such a trend (even when taking into account that middle granularity is higher as top models are still joining the game), which means that no strong conclusion should be made excluding older models, switching benchmarks frequently or similar actions, at most, old models may be left out of BAT, but other effects seem more pressing.
9.3 Benchmark used for visualizations
The benchmarks we used include: AGI Eval (Zhong2023AGIEvalAH), Alpaca (v2) (alpaca_eval), and its length-adjusted version (Dubois2024LengthControlledAA), HuggingFace OpenLLM Leaderboard (open-llm-leaderboard), MMLU (Hendrycks2020MeasuringMM), Chatbot-Arena and MTBench (Zheng2023JudgingLW), Big Bench Hard (suzgun2022challenging). ARC (clark2018think), HellaSwag (zellers2019hellaswag), TruthfulQA (lin2022truthfulqa), Winogrande (sakaguchi2019winogrande), EQ-Bench (v2) (paech2023eqbench).
NeurIPS Paper Checklist
1.
Claims
x
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
x
[…]
x
Justification: Sections 3,4,5
x
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
x
Question: Does the paper discuss the limitations of the work performed by the authors?
x
[…]
x
Justification: Limitations section
x
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory Assumptions and Proofs
x
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental Result Reproducibility
x
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
x
[…]
x
Justification: section 2
x
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
x
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
x
[…]
x
Justification: Upon acceptance, company does not allow to submit annonymized
x
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental Setting/Details
x
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
x
[…]
x
Justification: Section 2
x
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment Statistical Significance
x
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
x
[…]
x
Justification: Yes
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments Compute Resources
x
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code Of Ethics
x
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
x
[…]
x
Justification: Yes
x
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader Impacts
x
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
x
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
x
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
x
[…]
x
Justification: Section 2 and Appendix
x
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New Assets
x
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and Research with Human Subjects
x
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
x
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@ 6 BAT uses in Related Work
Some works survey and analyze a field by utilizing BAT techniques. Liu et al. (2021) check agreement across many QA datasets and conclude that since agreement is high, there is no need for more QA datasets. Sun et al. (2023) use correlations to show that Compositionality Benchmarks do not agree amongst themselves. They used Kendall-Tau and set as the high agreement threshold. Other works performed general efficient evaluation research and utilized BAT (Prabhu et al., 2024; Perlitz et al., 2023; Polo et al., 2024; Viswanathan et al., 2023). All of these works performed a thoughtful evaluation and large (reliable) rank correlation over all the models in the benchmarks. However, they did not consider the high correlations achieved in such settings (§3.2).
Other work relies on BAT to compare to a specific benchmark. Feng et al. (2024) automatically sample a small set of instructions as an efficient LLM benchmark, reducing human labor significantly. They show this still agrees with existing benchmarks. Similarly, Lei et al. (2023) and Viswanathan et al. (2023) both propose a synthetic benchmark as a proxy and show good agreement with the original benchmark, although they differ in their methodology. Chang et al. (2023) propose two benchmarks and use agreement to show that they capture the same phenomenon, and Mizrahi et al. (2023) test agreement within the same benchmarks using different prompts. Li et al. (2024b) validate a new benchmark with 6 models of 3 sizes 7B,13B,33B with agreement alpaca(v2) (Li et al., 2023). Yuan et al. (2024) and (Waldis et al., 2024) show divergent validity by comparing their benchmark to established ones, showing low BAT scores. Lastly, (Perlitz et al., 2023) compared efficient versions of the HELM benchmark to the full one.
@@ 7 Discussion and Conclusions
In this work, we shine a light on the lack of consistent BAT methodology. We analyze several BAT choices on a broad spectrum of benchmarks and assess their effect. Our analysis shows that different choices of (1) Models (2) Reference Benchmark(s), and (3) Thresholding scheme, can significantly alter BAT conclusions. Therefore, we advise a set of best practices and provide a Python package that aims to facilitate a consistent BAT process in the community. We also release the BenchBench-leaderboard, a benchmark that quantifies the agreement of a benchmark with an aggregate of existing benchmarks.
@@
It is not trivial how one should treat two benchmarks that are in high agreement with each other. If one is more convenient to run (e.g., doesn’t require costly metrics), then from a practical perspective, a user can simply choose it over the more expensive one. However, practitioners and researchers must not confuse high agreement with the notion that the benchmarks actually measure the exact same qualities. Among other things, this could lead to the erroneous conclusion that new benchmarks are no longer needed, impeding new benchmark development. The community must also discriminate between correlations of model abilities (strong models are strong at many tasks) and correlations of the benchmarks themselves (the benchmarks actually measure the same qualities).
@@
Another point concerns the role of BAT for benchmark retirement, i.e., at what point do we decide that an old benchmark is no longer relevant and should be discarded. Currently the issue of retirement is viewed mainly from the perspective of saturation, where the community stops using benchmarks on which all new models succeed.
However, another reason to retire benchmarks may be that the mixture of abilities models are expected to possess has shifted over time. In this scenario, BAT can reveal that a certain benchmark is no longer viable.
In conclusion, our study enhances the precision and reliability of Benchmark Agreement Testing by establishing best practices and introducing the BenchBench Python package and leaderboard. These contributions foster standardized evaluations, enabling more accurate comparisons across benchmarks and setting a new direction for computational linguistics research.
@@ 8 Limitations
We note that finding low agreement may indicate one of two issues, both of which have negative implications. These issues should be addressed or interpreted differently. One option is that the benchmark measures something different from what it is supposed to and is hence not valid. That is the more common interpretation and calls for changes. Another option might be that the benchmark is just not reliable, intuitively its ranking is unstable and did not converge. In such cases, even the same benchmark may not agree with itself given small changes (subsets, seeds etc.), this usually calls for evaluating on more examples (Choshen et al., 2024) or configuration (Bandel et al., 2024). There is a positive note to the same story, if a benchmark already shows a strong BAT in fine-grained evaluation (e.g., 5 models close to each other), it also means that it is quite reliable.
@@
In general, BAT needs a reference benchmark, or ideally multiple benchmarks that provide diverse measurements of the same construct. Still, choosing the right reference benchmarks might be tricky, and the results might be sensitive to this choice.